# Rethinking Bottlenecks in Safety Fine-Tuning of Vision Language Models

### v1

@@ Abstract
Large Vision-Language Models (VLMs) have achieved remarkable performance across a wide range of tasks. However, their deployment in safety-critical domains poses significant challenges. Existing safety fine-tuning methods, which focus on textual or multimodal content, fall short in addressing challenging cases or disrupt the balance between helpfulness and harmlessness. Our evaluation highlights a safety reasoning gap: these methods lack safety visual reasoning ability, leading to such bottlenecks. To address this limitation and enhance both visual perception and reasoning in safety-critical contexts, we propose a novel dataset that integrates multi-image inputs with safety Chain-of-Thought (CoT) labels as fine-grained reasoning logic to improve model performance. Specifically, we introduce the Multi-Image Safety (MIS) dataset, an instruction-following dataset tailored for multi-image safety scenarios, consisting of training and test splits. Our experiments demonstrate that fine-tuning InternVL2.5-8B with MIS significantly outperforms both powerful open-source models and API-based models in challenging multi-image tasks requiring safety-related visual reasoning. This approach not only delivers exceptional safety performance but also preserves general capabilities without any trade-offs. Specifically, fine-tuning with MIS increases average accuracy by 0.83% across five general benchmarks and reduces the Attack Success Rate (ASR) on multiple safety benchmarks by a large margin. Data and Models are released under: https://dripnowhy.github.io/MIS/ NOTE: This paper contains harmful images & text examples.
Large Vision Language Models, Multi-Image Safety, Dataset, Safety Fine-Tuning
@@
1 Introduction
Large Vision-Language models (VLMs) (Li et al., 2023; Liu et al., 2024b; Achiam et al., 2023; Team et al., 2024; Chen et al., 2024c) have emerged with exceptional visual and textual understanding capabilities, enabling them to perform excellently on multimodal tasks. With improved instruction-following abilities, Gong et al. (2023); Chen et al. (2024b); Liu et al. (2025); Zhang et al. (2024b) have focused on models’ ability to provide harmless responses when faced with image-text pairs containing unsafe elements. However, introduced visual information often bypasses the model’s safety mechanisms (Ding et al., 2024), posing great challenges to application in safety-critical tasks. To steer VLMs toward safer behaviors, researchers have introduced external safety feedback, such as Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022; Zhang et al., 2024b) and Supervised Fine-Tuning (SFT) (Zong et al., 2024; Li et al., 2024d), which effectively reduces the attack success rate against jailbreaks.
However, Guo et al. (2024b) highlights a major downside: these fine-tuning-based strategies often lead to over-conservatism. Even when faced with neutral or benign input, the fine-tuned models may still refuse to answer as shown in Figure 1 (b). Furthermore, our findings on more challenging safety-related tasks (Wang et al., 2024c; Zhou et al., 2024), where unsafe responses are triggered by neutral image-text pairs that tend to create unsafe intent, reveal a significant bottleneck in the safety capabilities of existing safeguarding methods. For example, as MSSBench (Zhou et al., 2024) shown in Figure 1 (a), (b), different images paired with benign text can convey the same intention, resulting in both safe and unsafe scenarios. The model must apply visual perception and reasoning to provide accurate responses in these different scenarios. However, the cases of Textual SFT and VLGuard in Figure 1 show that existing methods fail to generate appropriate responses. Our analysis reveals a safety visual reasoning gap: these methods lack the safety visual reasoning capabilities, resulting in failures to balance the helpfulness-harmlessness trade-off and to address complex safety scenarios that cannot be addressed by safety mechanisms relying solely on basic visual perception.
To bridge this gap and enhance both safety-related visual perception and visual reasoning in safety scenarios, we introduce the Multi-Image Safety (MIS) dataset, which contains both training and test splits. We developed an automated data generation framework using LLMs, VLMs, and Text-to-Image (T2I) models, and deployed a multi-expert system composed of human experts and GPT-4o (Achiam et al., 2023) to clean and classify the generated data. As shown in Figure 1 (c), the MIS samples utilize text instructions to combine two images, introducing unsafe intent that the model must interpret and address. The model is required to apply both visual perception and reasoning to generate a safe response. For the MIS training set, we generate responses using InternVL with safety CoT prompt. By fine-tuning with it, the model’s safety capabilities on challenging multimodal safety tasks can be largely improved without any trade-off in general capabilities. Our main contributions and findings are summarized as follows:
@@ 2 Bottlenecks in Safety Fine-Tuning Vision Language Models
Fine-tuning-based methods face significant bottlenecks in the safety alignment of Vision Language Models (VLMs). Hu et al. (2024) highlighted that safety capabilities acquired through Textual Supervised Fine-Tuning often fail to generalize to the visual domain, suggesting that Multimodal SFT could be a more viable alternative. However, Ding et al. (2024); Guo et al. (2024b) observed that existing Multimodal SFT methods tend to exhibit over-prudence, frequently refusing to respond even to safe image-text inputs. Furthermore, with the emergence of challenging safety-related tasks, such as those that trigger harmful responses from models using safe images combined with safe text (Wang et al., 2024c; Zhou et al., 2024), which require a certain degree of reasoning to identify the underlying unsafe intent, we have found that existing SFT methods are insufficient to provide effective defenses. Our experiments and analysis indicate that the cause of safety bottlenecks can primarily be attributed to two factors: (i) composition of SFT inputs and (ii) construction method of SFT labels.
2.1 Experimental Setups
Model and Baselines.
We use LLaVA-v1.5-13B (Liu et al., 2024a), Qwen2-VL-7B (Wang et al., 2024b), and InternVL2.5-8B (Chen et al., 2024c) as our base models. LLaVA is one of the most commonly used VLM, while the other two are recent models with strong general capabilities. For Textual SFT, we follow the setup in Hu et al. (2024) and fine-tune with SafeRLHF (Ji et al., 2024). For Multimodal SFT, we use VLGuard (Zong et al., 2024), which includes 2k unsafe samples and 1k benign inputs. As we cannot access the training data of recent SOTA VLMs, we apply Posthoc method (VLGuard-P) on these models. For LLaVA, we test the Mixed method VLGuard-M and VLGuard-P. To study the impact of SFT labels on model performance, we reconstructed the labels of VLGuard, referring to this method as VLGuard-R. More details are shown in Section A.1.
Benchmarks.
For general tasks, we select the following benchmarks: MMStar (Chen et al., 2024a) for Single-Image (SI), MMMU (Yue et al., 2024) for Single-Image and Multi-Image (SI&MI), and MMT-Bench-MI (Ying et al., 2024) for Multi-Image (MI). For safety-related tasks, we primarily use two challenging datasets: FigStep (Gong et al., 2023), where harmful information is converted from text to image via OCR with benign text input, and MSSBench (Zhou et al., 2024), which constructs situational safety scenarios using different images, posing great challenges for VLM safety. We provide more information in Section A.2.
Metrics.
We use VLMEvalKit (Duan et al., 2024) and lmms-eval (Zhang et al., 2024a) to evaluate the general capabilities on benchmarks and report the relevant metrics. For safety-related tasks, we report the Attack Success Rate (ASR) on FigStep using the SOTA safety assessment model LlamaGuard3-8B (Inan et al., 2023). On MSSBench, we follow settings in their paper, using GPT-4o (Achiam et al., 2023) to classify responses as safe or unsafe, and then calculate their accuracy across different situations.
Table 1: Comparison of different SFT methods on three base VLMs: LLaVA-v1.5-13B, Qwen2-VL-7B, and InternVL2.5-7b across general and safety tasks. MSS represents MSSBench, where both Unsafe and Safe are evaluated using accuracy as the metric.
MMStar SI
MMMU SI&MI
MMT MI
FigStep
MSS
@@ 2.1 Experimental Setups
2.2 Findings and Disccusions
@@
Discussion 1: What Causes Collapse of General Capabilities in Multimodal SFT?
We examined the cases where VLGuard failed and found that most responses began with “I’m sorry”. We hypothesize that this is due to the labels of the unsafe sample in VLGuard SFT, which predominantly consist of simple rejection responses starting with “I’m sorry”. Fine-tuning models on such data leads to over-prudence on visual features, causing the model to reject benign visual inputs. To validate our hypothesis, we conduct further experiments on the safe situations in MSSBench. For the same safe instruction, we use three different input formats: (i) paired with the corresponding safe image from MSSBench, (ii) paired with a white image, and (iii) text input only. The Reject Rate (RR) for each format is reported in Table 2. Results indicate that the VLGuard-P fine-tuned model exhibits severe over-prudence in the visual domain. For safe instructions, even when paired with a meaningless white image, the model generates nearly 50% rejection responses. In contrast, with text-only input, the model’s RR is significantly lower. This validates vanilla Multimodal SFT leads to excessive conservatism in the visual domain.
@@
Safe Instruction + Related Image 32.57 64.37
Safe Instruction + White Image 42.15 59.33
@@
Safe Instruction + Related Image 42.50 63.33
Safe Instruction + White Image 48.65 62.86
@@
The excessive simple rejection response in Multimodal SFT labels leads to over-prudence, raising a direct question: Can this phenomenon be mitigated by constructing better SFT labels? Guo et al. (2024a); Xu et al. (2024) demonstrated that the construction of chain-of-thought (CoT) templates for SFT labels improves the visual understanding ability of VLMs, which inspires us to explore whether models could provide safe responses by performing simple reasoning on image-text pairs, instead of refusing. Thus, for the 2k unsafe data in VLGuard, we prompt InternVL2.5-78B (Chen et al., 2024c) to perform simple reasoning to analyze why the input is unsafe and then provide a safe response. The prompt is provided in Section A.2. As shown in Table 1, the model fine-tuned with VLGuard-R demonstrates general performance comparable to the base VLMs and significantly outperforms VLGuard-P.
@@
Interestingly, although Textual SFT and VLGuard demonstrate decent performance on FigStep, these methods perform poorly on the more challenging MSSBench. As shown in the results in Table 1, both VLGuard-M and Textual SFT fail to provide warning or advisory safety responses in unsafe situations, resulting in unsafe accuracy similar to the base model, both being less than 10%. Regarding VLGuard-P, although it achieves an approximate accuracy 60% in unsafe situations, the results in Table 2 indicate that this is not due to its ability to recognize harmful intent. Instead, it tends to give rejection responses for most images, even the white image, resulting in similar accuracy in safe and unsafe situations.
@@
Current challenging safety-related tasks are designed to trigger risky responses from models by using safe text and safe image inputs (Wang et al., 2024c). The difficulty of MSSBench (Zhou et al., 2024) lies in that the same safe query, when paired with different safe images, can build both safe and unsafe situations. The model is expected to provide helpful responses in safe situations and issue warnings or highlight potential safety risks in unsafe situations. We believe that addressing such challenges requires models to possess both visual perception and reasoning capabilities. Specifically, the model must recognize visual content and reason about it based on the text query to provide contextually accurate responses. Textual SFT, relying solely on text inputs, cannot enhance visual capabilities. VLGuard, which uses single-image and text inputs, tends to match specific visual elements as safe or unsafe, lacking the capability to reason about potential harmful intent. Even with VLGuard-R, despite construct labels with some reasoning logic, the input data primarily consists of simple unsafe elements. As shown in Table 1, the performance on challenging tasks remains relatively limited.
@@ 3 Multi-Image Safety Fine-tuning
To overcome the previously mentioned bottlenecks, in Sec. 3.2, we construct the first Multi-Image Safety dataset (MIS), designed to enhance model safety-related ability on both visual perception and reasoning. Building on the MIS training set, we propose a fine-tuning method to gain safety-related visual perception and reasoning abilities from multi-image: Multi-Image Reasoning Safety (MIRage) in Sec. 3.3.
@@ 3.1 Overview of MIS
Our MIS containing 4k training sampples and 2185 testing samples, designed to improve and evaluate model safety-related ability on both visual perception and reasoning. The MIS test set consists of three subsets: MIS-easy, MIS-hard, and MIS-real. Each sample contains a text query and two images, with images in MIS-easy and MIS-hard generated using a Text-to-Image (T2I) model, while those in MIS-real are retrieved from existing image datasets. In this multi-image setting, harmful intent arises from the relationships between images, while the text queries remain neutral as shown in Figure 3. We differentiate MIS-easy and MIS-hard based on whether the images contain explicit unsafe elements. For example, the hard sample related to privacy in Figure 3 involves a camera and a bedroom in separate images. The risk of illegal surveillance arises only when these benign images are combined with the instruction.
Taxonomy.
Regarding our data construction based on harmful text and images from existing benchmarks, following previous work on multimodal and LLM safety (Hu et al., 2024; Li et al., 2024c; Liu et al., 2025), we cover a total of 6 key unsafe categories and 12 subcategories. The detail count and the ratio of MIS test set are provided in Table 3. You can find more examples in Appendix D.
@@ 3.2 Construction Pipeline
We present the MIS dataset construction pipeline in Figure 2, which generates high-quality multi-image-text pairs with safety risks through four key steps. In Step 1, Qwen2.5-72B (Yang et al., 2024) and InternVL2.5-78B (Chen et al., 2024c) are prompted to extract unsafe elements from the text and images of existing safety-related benchmarks. In Step 2, a few-shot prompt is designed to guide Qwen2.5-72B in generating harmful queries involving two objects based on unsafe elements. The objects in these harmful queries are then replaced with phrases like “xxx in the image”. The refined queries are subsequently detoxified to produce the final text instructions. As a result, we obtain two objects and a safe text instruction for each sample. Step 3 involves an auto-refinement image generation process using Stable Diffusion 3.5 Large (Esser et al., 2024). In the first round, images are generated based on the objects identified in Step 2. InternVL2.5-78B then refines the T2I prompts using contextual information from Step 2 to improve the alignment between images and text. By leveraging these refined prompts, the consistency between the second round of generated images and the text instructions is significantly enhanced. Finally, in Step 4, human experts and GPT-4o (Achiam et al., 2023) filter and classify the generated image-text pairs. Pairs with dangerous intent in text instructions are assigned to the training set, neutral text paired with explicit harmful elements in images is categorized as MIS-easy, and neutral text-image pairs with no harmful elements are classified as MIS-hard. This comprehensive process ensures the creation of a high-quality dataset for multi-image safety evaluation.
@@
As mentioned in Sec. 3.2, the training set of MIS is filtered by assessing the safety risks in the text instructions. In the visual domain, this includes not only images with potentially unsafe risk but also challenging cases where the completely unsafe intent arises from the relationship between two safe images. The multi-image inputs allow us to easily create such samples, in contrast to previous SFT methods that were solely built on simple unsafe elements. Additionally, In Sec. 2.2, we verified that incorporating visual reasoning logic into the SFT labels improves performance on both general and challenging safety-related tasks. Considering the multi-image scenario we constructed, the ideal model response should first understand the content in the images, then reason about the unsafe intent based on the relationship between the two images and the text, and finally provide a safe response. Therefore, we design a safety CoT prompt to guide InternVL2.5-78B in generating safe responses that contain simple visual perception and visual reasoning logic. The prompt is provided in Section E.2.
Finally, we get 4k samples for the training set, 1675 for MIS-easy, and 510 for MIS-hard. In addition, 100 test samples are used for real image retrieval. Specifically, 200 images from the LAION-2B-en-research-safe dataset (Schuhmann et al., 2022) are matched with objects and paired with text instructions to form the MIS-real subset. We provide a detailed illustration of the entire pipeline in Section C.1.

### v3

@@ Abstract
Large Vision-Language Models (VLMs) have achieved remarkable performance across a wide range of tasks. However, their deployment in safety-critical domains poses significant challenges. Existing safety fine-tuning methods, which focus on textual or multimodal content, fall short in addressing challenging cases or disrupt the balance between helpfulness and harmlessness. Our evaluation highlights a safety reasoning gap: these methods lack safety visual reasoning ability, leading to such bottlenecks. To address this limitation and enhance both visual perception and reasoning in safety-critical contexts, we propose a novel dataset that integrates multi-image inputs with safety Chain-of-Thought (CoT) labels as fine-grained reasoning logic to improve model performance. Specifically, we introduce the Multi-Image Safety (MIS) dataset, an instruction-following dataset tailored for multi-image safety scenarios, consisting of training and test splits. Our experiments demonstrate that fine-tuning InternVL2.5-8B with MIS significantly outperforms both powerful open-source models and API-based models in challenging multi-image tasks requiring safety-related visual reasoning. This approach not only delivers exceptional safety performance but also preserves general capabilities without any trade-offs. Specifically, fine-tuning with MIS increases average accuracy by 0.83% across five general benchmarks and reduces the Attack Success Rate (ASR) on multiple safety benchmarks by a large margin.
NOTE: This paper contains harmful images & text examples.
1 Introduction
Large Vision-Language models (VLMs) (Liu et al., 2024b; Achiam et al., 2023; Team et al., 2024; Chen et al., 2024c) have emerged with exceptional visual and textual understanding capabilities, enabling them to perform excellently on multimodal tasks. With improved instruction-following abilities, Gong et al. (2023); Chen et al. (2024b); Liu et al. (2025) have focused on models’ ability to provide harmless responses when faced with image-text pairs containing unsafe elements. However, introduced visual information often bypasses the model’s safety mechanisms (Ding et al., 2024), posing great challenges to application in safety-critical tasks. To steer VLMs toward safer behaviors, researchers have introduced external safety feedback, such as Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022; Zhang et al., 2024b) and Supervised Fine-Tuning (SFT) (Zong et al., 2024; Li et al., 2024d), which effectively reduces the attack success rate against jailbreaks.
However, Guo et al. (2024b) highlights a major downside: these fine-tuning-based strategies often lead to over-conservatism. Even when faced with neutral or benign input, the fine-tuned models refuse to answer as shown in Fig. 1 (b). Furthermore, our findings on more challenging safety tasks (Wang et al., 2024c; Zhou et al., 2024), where unsafe responses are triggered by neutral image-text pairs that create unsafe intent, reveal a bottleneck in the safety capabilities of existing safeguarding methods. For example, as MSSBench (Zhou et al., 2024) shown in Fig. 1 (a), (b), different images paired with benign text can convey the same intention, resulting in both safe and unsafe scenarios. The model must apply visual perception and reasoning to provide accurate responses in these different scenarios. However, the cases of Textual SFT and VLGuard in Fig. 1 show that existing methods fail to generate appropriate responses. Our analysis reveals a safety visual reasoning gap: existing methods lack the ability to reason about safety in visual contexts, leading to failures in balancing helpfulness and harmlessness and in handling complex scenarios beyond basic visual perception.
@@ 1 Introduction
To bridge this gap and enhance both safety-related visual perception and visual reasoning in safety scenarios, we introduce the Multi-Image Safety (MIS) dataset, which contains both training and test splits. We develop an automated data generation framework using LLMs, VLMs, and Text-to-Image models, and deploy a multi-expert system composed of human experts and GPT-4o to clean and classify the generated data. As shown in Fig. 1 (c), the MIS example utilize text instructions to combine two images, introducing unsafe intent that the model must interpret and address. The model is required to apply both visual perception and reasoning to generate a safe response. For the MIS training set, we generate responses using InternVL with safety CoT prompt. By fine-tuning with it, the model’s safety performance on challenging safety tasks can be largely improved without any trade-off in general capabilities. Our main contributions and findings are summarized as follows:
@@ 2 Bottlenecks in Safety Fine-Tuning Vision Language Models
Table 1: Comparison of different SFT methods on three base VLMs: LLaVA-v1.5-13B, Qwen2-VL-7B, and InternVL2.5-8b across general and safety tasks. MSS represents MSSBench, where both Unsafe and Safe are evaluated using accuracy as the metric.
MMStar SI MMMU SI&MI MMT-Bench MI FigStep MSS Safe MSS Unsafe
@@
Fine-tuning-based methods face significant bottlenecks in safeguarding VLMs. Hu et al. (2024) highlighted that safety capabilities acquired through Textual Supervised Fine-Tuning often fail to generalize to the visual domain, suggesting that Multimodal SFT could be a more viable alternative. However, Ding et al. (2024); Guo et al. (2024b) observed that existing Multimodal SFT methods tend to exhibit over-prudence, frequently refusing to respond even to safe image-text inputs. Furthermore, with the emergence of challenging safety-related tasks, such as those that trigger harmful responses from models using safe images combined with safe text (Wang et al., 2024c; Zhou et al., 2024), which require a certain degree of reasoning to identify the underlying unsafe intent, we have found that existing SFT methods are insufficient to provide effective defenses. Our experiments and analysis indicate that the cause of safety bottlenecks is likely to be attributed to two factors: (i) composition of SFT inputs and (ii) construction method of SFT labels.
2.1 Experimental Setups
Model and Baselines.
We use LLaVA-v1.5-13B (Liu et al., 2024a), Qwen2-VL-7B (Wang et al., 2024b), and InternVL2.5-8B (Chen et al., 2024c) as our base models. To demonstrate the completeness of our analysis, we also include a discussion of the reasoning model MiMo-VL-7B-RL in Appendix C.1. LLaVA is one of the most commonly used VLM, while the other two are recent models with strong general capabilities. For Textual SFT, we follow the setup in (Hu et al., 2024) and fine-tune with SafeRLHF (Ji et al., 2024). For Multimodal SFT, we use VLGuard (Zong et al., 2024), which includes 2k unsafe samples and 1k benign inputs. As we cannot access the training data of recent SOTA VLMs, we apply Posthoc method (VLGuard-P) on these models. For LLaVA, we test the Mixed method VLGuard-M, VLGuard-P, and Textual SFT. To study the impact of SFT labels on model performance, we reconstructed the labels of VLGuard, referring to this method as VLGuard-R. More details are shown in Appendix B.2.
Benchmarks.
For general tasks, we select the following benchmarks: MMStar (Chen et al., 2024a) for Single-Image (SI), MMMU (Yue et al., 2024) for Single-Image and Multi-Image (SI&MI), and MMT-Bench-MI (Ying et al., 2024) for Multi-Image (MI). For safety-related tasks, we primarily use two challenging datasets: FigStep (Gong et al., 2023), where harmful information is converted from text to image via OCR with benign text input, and MSSBench (Zhou et al., 2024), which constructs situational safety scenarios using different images, posing great challenges for VLM safety. We provide more information in Appendix B.3.
Metrics.
We use VLMEvalKit (Duan et al., 2024) to evaluate the general capabilities on benchmarks and report the relevant metrics. For safety-related tasks, we report the Attack Success Rate (ASR) on FigStep using the SOTA safety assessment model LlamaGuard3-8B (Inan et al., 2023). On MSSBench, we follow settings in their paper, using GPT-4o (Achiam et al., 2023) to classify responses as safe or unsafe, and then calculate their accuracy across different situations.
2.2 Findings and Discussions
@@
Discussion 1: What Causes the Collapse of General Capabilities in Multimodal SFT?
We examined the cases where VLGuard failed and found that most responses began with ”I’m sorry”. We hypothesize that this is due to the labels of the unsafe sample in VLGuard SFT, which predominantly consist of simple rejection responses starting with ”I’m sorry”. Fine-tuning models on such data leads to over-prudence on visual features, causing the model to reject benign visual inputs. To validate our hypothesis, we conduct further experiments on the safe situations in MSSBench.
@@
Safe Instr. + Related Image 32.57 64.37
Safe Instr. + White Image 42.15 59.33
@@
Safe Instr. + Related Image 42.50 63.33
Safe Instr. + White Image 48.65 62.86
@@
For the same safe instruction, we use three different input formats: (i) paired with the corresponding safe image from MSSBench, (ii) paired with a white image, and (iii) text input only. The Reject Rate (RR) for each format is reported in Table 2. Results indicate that the VLGuard-P fine-tuned model exhibits severe over-prudence in the visual domain. For safe instructions, even when paired with a meaningless white image, the model generates nearly 50% rejection responses. In contrast, with text-only input, the model’s RR is significantly lower. This validates vanilla Multimodal SFT leads to excessive conservatism in the visual domain.
@@
The excessive simple rejection response in Multimodal SFT labels leads to over-prudence, raising a direct question: Can this phenomenon be mitigated by constructing better SFT labels? Guo et al. (2024a); Xu et al. (2024) demonstrated that the construction of chain-of-thought (CoT) templates for SFT labels improves the visual understanding ability of VLMs, which inspires us to explore whether models could provide safe responses by performing simple reasoning on image-text pairs, instead of refusing. Thus, for the 2k unsafe data in VLGuard, we prompt InternVL2.5-78B (Chen et al., 2024c) to perform simple reasoning to analyze why the input is unsafe and then provide a safe response. The prompt is provided in Appendix B.3. As shown in Table 1, the model fine-tuned with VLGuard-R demonstrates general performance more comparable to the base VLMs and outperforms VLGuard-P.
@@
Interestingly, although Textual SFT and VLGuard demonstrate decent performance on FigStep, these methods perform poorly on the more challenging MSSBench. As shown in the results in Table 1, both VLGuard-M and Textual SFT fail to provide warning or advisory safety responses in unsafe situations, resulting in unsafe accuracy similar to the base model, both being less than 10%. Regarding VLGuard-P, although it achieves an approximate accuracy 60% in unsafe situations, the results in Table 2 indicate that this is not due to its ability to recognize harmful intent. Instead, it tends to give rejection for most images, even the white image, resulting in similar accuracy in safe and unsafe situations.
@@
Current challenging safety-related tasks are designed to trigger risky responses from models by using safe text and safe image inputs (Wang et al., 2024c). The difficulty of MSSBench (Zhou et al., 2024) lies in that the same safe query, when paired with different safe images, can build both safe and unsafe situations. The model is expected to provide helpful responses in safe situations and issue warnings or highlight potential safety risks in unsafe situations. We believe that addressing such challenges requires models to possess both visual perception and reasoning capabilities. Specifically, the model must recognize visual content and reason about it based on the text query to provide contextually accurate responses. Textual SFT, relying solely on text inputs, cannot enhance visual capabilities. VLGuard, which uses single-image and text inputs, tends to match specific visual elements as safe or unsafe, lacking the capability to reason about potentially harmful intent. Even with VLGuard-R, despite constructing labels with some reasoning logic, the input data primarily consists of simple unsafe elements. As shown in Table 1, the performance on challenging tasks remains relatively limited.
@@ 3 Multi-Image Safety Fine-tuning
To address the aforementioned challenges, we introduce the first Multi-Image Safety dataset (MIS) in Sec. 3.2, designed to improve safety-related visual perception and reasoning. Based on the MIS training set, we further propose MIRage (Multi-Image Reasoning Safety), a fine-tuning approach for enhancing multi-image safety capabilities.
@@ 3.1 Overview of MIS
Our MIS dataset, comprising 4k training and 2185 testing samples, is designed to evaluate safety in both visual perception and reasoning. Each test sample includes a neutral text query and two images, with images in MIS-easy and MIS-hard generated by a T2I model, and those in MIS-real drawn from existing datasets. Harmful intent arises from image–image relationships rather than the text, as illustrated in Fig. 2. We distinguish easy and hard cases by whether the images contain explicit unsafe elements; for instance, combining benign images of a camera and a bedroom can imply illegal surveillance (Fig. 2). Following prior work on multimodal and LLM safety (Hu et al., 2024; Li et al., 2024c; Liu et al., 2025), MIS covers 6 unsafe categories and 12 subcategories, with distributions detailed in Table 3 and more examples in Appendix E.
@@ 3.2 Construction Pipeline
We present the MIS dataset construction pipeline in Fig. 3, which generates high-quality multi-image-text pairs with safety risks through four key steps. In Step 1, Qwen2.5-72B (Yang et al., 2024) and InternVL2.5-78B (Chen et al., 2024c) are prompted to extract unsafe elements from the text and images of existing safety-related benchmarks. In Step 2, a few-shot prompt is designed to guide Qwen2.5-72B in generating harmful queries involving two objects based on unsafe elements. The objects in these harmful queries are then replaced with phrases like ”xxx in the image”. The refined queries are subsequently detoxified to produce the final text instructions. As a result, we obtain two objects and a safe text instruction for each sample. Step 3 involves an auto-refinement image generation process using Stable Diffusion 3.5 Large (Esser et al., 2024). In the first round, images are generated based on the objects identified in Step 2. InternVL2.5-78B then refines the T2I prompts using contextual information from Step 2 to improve the alignment between images and text. By leveraging these refined prompts, the consistency between the second round of generated images and the text instructions is significantly enhanced. Finally, in Step 4, human experts and GPT-4o filter and classify the generated image-text pairs. Pairs with dangerous intent in text instructions are assigned to the training set, neutral text paired with explicit harmful elements in images is categorized as easy, and neutral text-image pairs with no harmful elements are classified as hard. This comprehensive process ensures the creation of a high-quality dataset for multi-image safety evaluation.
@@
As described in Sec. 3.2, the MIS training set is filtered based on safety risks in text instructions. In the visual domain, this includes not only unsafe images but also cases where unsafe intent arises from the interaction of two otherwise safe images, a construction enabled by multi-image inputs and absent in prior SFT approaches. In Sec. 2.2, we show that adding visual reasoning to SFT labels improves performance on both general and safety-critical tasks. In the multi-image setting, an ideal model should comprehend the visual content, reason about unsafe intent from image–text relationships, and then provide a safe response. To support this, we design a safety CoT prompt that guides InternVL2.5-78B to generate responses integrating visual perception and reasoning.
Finally, we get 4k samples for the training set, 1675 for MIS-easy, and 510 for MIS-hard. In addition, 100 samples are used for real image retrieval. Specifically, 200 images from the LAION-2B-en-research-safe dataset (Schuhmann et al., 2022) are matched with objects and paired with text instructions to form the MIS-real. We provide a detailed illustration of the pipeline in Appendix D.1.
MIRage.
Similar to prior SFT methods, we add 500 general QA samples from M4-Instruct (Li et al., 2024b) to preserve instruction-following ability. In the final 4.5k training set, only 11% are general samples, far lower than the 33% used in Textual SFT and VLGuard (Zong et al., 2024). The full training process is detailed in Appendix B.2.
4 Experiments
Training Details.
Our main experiments focus on applying MIRage to InternVL2.5-8B (Chen et al., 2024c), with results on additional models presented in Appendix C. More training details also are provided in Appendix B.2.
Baselines.
We evaluate 14 VLMs with robust multi-image understanding capabilities, including 11 open-sourced models and 2 closed-source API models. Additionally, we also evaluate some safety fine-tuning baselines, including Textual SFT and VLGuard-R on 4 open-sourced VLMs. A detailed baseline introduction is provided in Appendix B.3.
Evaluation on MIS Test Set.
We use GPT-4o (Achiam et al., 2023) to classify all responses into four categories: Unsafe, for responses that present safety risks based on harmful instructions; Safe with Reasoning, for answers that identify content in the images and logically deduce potential harmful intent with a warning; Safe with Refusal, for brief refusal responses; and Hallucination, for responses that are irrelevant or incomplete due to the model’s failure to understand the samples. Based on these four categories, we use Attack Success Rate (ASR), Reasoning Success Rate (RSR), Refusal Rate (RR), and Hallucination Rate (HR) as metrics to evaluate the model’s safety capabilities.
Evaluation Benchmarks on MIRage.
We evaluate our MIRage from two perspectives: Harmlessness and Helpfulness. Building on the settings from Sec. 2.1, we conduct additional experiments on general benchmarks such as MuirBench (Wang et al., 2024a), and Q-Bench (Wu et al., 2023). Furthermore, we introduce more safety tasks, including MM-SafetyBench (Liu et al., 2025), Jailbreak-V mini-set (Luo et al., 2024), and the challenging SIUO task (Wang et al., 2024c). We also evaluate MIRage on MOSSBench (Li et al., 2024e), which is provided in Appendix C.2. Detailed settings and benchmark description can be found in Appendix B.3.
4.1 Main Result